For csc^2 a- 1/cotA cscA--what is the simplest equivalent trigonometric expression
Answer by venugopalramana(3286)

For csc^2 a- 1/cotA cscA--what is the simplest equivalent trigonometric expression
HOPE YOU MEAN
(csc^2 a- 1)/cotA cscA
=COT^2(A)/COT(A)COSEC(A)=COT(A)/COSEC(A)=(COS(A)/SIN(A))/(1/SIN(A))=COS(A)